# Linz, Austria

### v1

@@ Abstract
We present f4ncgb, a new open-source C++ library for Gröbner basis computations in free algebras, which transfers recent advancements in commutative Gröbner basis software to the noncommutative setting. As our experiments show, f4ncgb establishes a new state-of-the-art for noncommutative Gröbner basis computations. We also discuss implementation details and design choices.
@@ 1 Introduction
In this work, we present f4ncgb, a new open-source C++ library for Gröbner basis computations in free algebras, distributed under the license MIT and available at https://gitlab.sai.jku.at/f4ncgb/f4ncgb. The library supports noncommutative polynomial systems with coefficients in prime fields of characteristic $2^{31}$ and in the field of rational numbers. To the best of our knowledge, f4ncgb is the first open-source tool to port the newest advancements in Gröbner basis computations from the commutative setting to the noncommutative world. Our experiments in Section 8 show that f4ncgb establishes a new state-of-the-art for noncommutative Gröbner basis computations among freely available tools.
The foundation of our implementation are new data structures for representing and storing noncommutative monomials and polynomials. We abstract these data types into a store structure, using a contiguous memory area for storing elements and a hash map for efficient lookup of known entries. We describe this structure in more detail in Section 5.
@@
To support this efficiently, we have implemented a specialized Gaussian elimination algorithm tailored for sparse matrices, based on ideas from the commutative case [33]. Our implementation makes use of modern multi-core hardware by parallelizing matrix reductions, resulting in significant overall speedups. For inputs with rational coefficients, we employ a multi-modular strategy: computations are carried out modulo different primes, followed by a reconstruction step to recover the rational solution. We provide an overview of the noncommutative F4 algorithm in Section 4 and discuss our linear algebra implementation in Section 6.
In addition to computing Gröbner bases, f4ncgb offers proof logging capabilities in the form of computing cofactor representations: for each polynomial in the output Gröbner basis, a representation in terms of the input can be computed, see Section 7. These representations are not only of interest for verifying the correctness of the output, but are also useful in applications [22].
@@ 2 Related Software
Notable examples of tools relying on Buchberger’s algorithm include the Mathematica library NCAlgebra [17], the ApCoCoA package NCPoly [28], and the GAP package GBNP [11]. Other tools that also fall into this category are the Standard Lisp implementation bergman [3] or the C++ tool Opal [16], which both seem no longer actively maintained.
An important exception is the subscription-based computer algebra system Magma [8], which is one of the few tools to implement a noncommutative version of the F4 algorithm. In addition, Magma also features dedicated data structures and highly optimized subroutines (e.g., for the linear algebra), making it arguably the state-of-the-art for computing noncommutative Gröbner bases. However, its significant drawback is that it is not freely accessible, limiting its availability to the community.
To the best of our knowledge, the only freely available and open-source implementation of the noncommutative F4 algorithm is the SageMath package operator_gb [5], developed by one of the authors. While operator_gb already implements the noncommutative F4 algorithm, it still relies on relatively simple data structures for representing and manipulating noncommutative objects. Additionally, the package is primarily written in interpreted Python, with selected components (such as the linear algebra) in compiled Cython, which limits performance. As a result, it cannot match the efficiency of the highly optimized and fully compiled C code in Magma. The new package f4ncgb can be viewed as a significant advancement of operator_gb, offering both improved algorithmic design and a more efficient implementation.
@@ 3 Preliminaries – Noncommutative Gröbner Bases
For a set of polynomials $F\subseteq K\langle X\rangle$, we denote by $(F)$ the (two-sided) ideal generated by $F$, consisting of all two-sided linear combinations of the elements in $F$ with polynomials as coefficients, i.e.,
$(F)=\left\{\sum_{i=1}^{k}p_{i}f_{i}q_{i}\mid f_{i}\in F,\;p_{i},q_{i}\in K\langle X\rangle,\;k\in\mathbb{N}\right\}.$
A set $F$ is a generating set of an ideal $I\subseteq K\langle X\rangle$ if $I=(F)$. Recall that if $|X|>1$, then there exist ideals in $K\langle X\rangle$ that do not possess a finite generating set. A classical example is the ideal $(\{xy^{n}x\mid n\in\mathbb{N}\})\subseteq K\langle x,y\rangle$.
@@
A very general class of monomial orderings are weighted block orderings of the following form: Fix $d\in\mathbb{Z}_{>0}$ and assign to each variable $x_{i}\in X$ a nonzero weight vector $w_{i}\in\mathbb{R}_{\geq 0}^{d}$. Define the weight $w(m)$ of a monomial $m=x_{i_{1}}\dots x_{i_{k}}\in\langle X\rangle$ as $w(m)=\sum_{j=1}^{k}w_{i_{j}}\in\mathbb{R}_{\geq 0}^{d}$. The weighted block ordering $\prec_{\omega}$ w.r.t. the weight matrix $\omega=(w_{1},\dots,w_{n})$ is given by:
@@
At the moment, our implementation supports weighted block orderings where each weight vector is a unit vector $(0,\dots,0,1,0,\dots,0)$. This is sufficient to represent classical orderings such as degree-lexicographic orderings. It also enables the use of elimination orderings [20, Def. 2.4.15], which are crucial, e.g., for computing elimination ideals [20, Thm. 2.4.43].
@@
We note that all standard characterizations of Gröbner bases from the commutative setting carry over naturally to the noncommutative case; see, e.g., [20, Thm. 2.4.37]. In particular, Buchberger’s criterion [13, Thm. 2.6.6], which provides an effective test for commutative Gröbner bases, has a noncommutative analogue in Bergman’s diamond lemma [4, Thm. 1.2].
@@
Analogous to Buchberger’s criterion, Bergman’s diamond lemma is based on a notion of (noncommutative) polynomial reduction, which is (basically) identical to the reduction used in the commutative case: $f$ can be reduced by $g$ if $\operatorname{lm}(g)$ divides a monomial in $f$ and the result of the reduction is $f-\lambda vgw$ with $\lambda\in K$, $v,w\in\langle X\rangle$ chosen so that the targeted monomial cancels. For a formal definition, see, e.g., [20, Sec. 2.4.2].
@@
Theorem 3.1 provides an effective criterion to test whether a given set forms a Gröbner basis. It also directly gives rise to a completion procedure for constructing a Gröbner basis from an arbitrary generating set – a noncommutative analogue of Buchberger’s algorithm; see [20, Alg. 1] for a detailed description.
@@ 4 Noncommutative F4 Algorithm
Our library f4ncgb implements the noncommutative F4 algorithm for computing Gröbner bases in free algebras. Below, we provide a brief overview of the algorithm; for a detailed exposition, see [39, Sec. 5.4]. For those familiar with the commutative algorithm, we note that the noncommutative version follows the same high-level structure.
After initializing the prospective Gröbner basis with the input polynomials and computing all ambiguities between them, the F4 algorithm proceeds with the following main steps:
@@
    Basis update: Translate the rows of the reduced matrix back into polynomials and add all those polynomials to the basis whose leading monomial is not divisible by an element of the current basis. Then, form all new ambiguities. Optionally, one can apply the noncommutative variant of the Gebauer-Möller criteria [39, Thm. 4.2.22] to discard redundant ambiguities. If ambiguities remain, continue with the next iteration.
@@ 5 Basic Data Structures – Monomials and Polynomials
To store monomials (and also polynomials), we designed a common base store structure that is extended into monomial store and polynomial store structures using the Curiously Recurring Template Pattern [12]. The store implements a bump allocator combined with an open-addressing hash map (using Boost’s unordered_flat_map) for efficient lookups of known entries.
@@
Similarly, we also maintain a global polynomial store that contains all polynomials generated during a Gröbner basis computation. Each polynomial entry consists of metadata and a list of its monomials. The monomials are represented by their index in the monomial store and are ordered in descending order according to the monomial ordering, making the leading monomial easily accessible as the first entry. The metadata consists of the polynomial’s length and of an index into a global coefficient array that stores the coefficients of all polynomials. We do not store the coefficients directly in the store as typically many polynomials share the same coefficients (because they are monomial multiplies of each other). In this way, we can store the coefficients just once and reference them from multiple polynomials. Rational coefficients are stored with arbitrary-precision. Figure 1 shows an example of a monomial and polynomial store.
@@ 6.1 Gaussian Elimination in Prime Fields
The forward elimination phase begins with a preprocessing step in which we permute the rows of $A$ such that, in each row, the first nonzero entry appears directly below or to the left of the first nonzero entry in the row above. This reordering yields a matrix with an approximately triangular structure, as illustrated in Figure 2. This preprocessing step is crucial for the efficiency of the algorithm.
[…]
[width=]gauss_elim
@@
Once the buffer is fully reduced, we copy its content back into the sparse row $r$. If the resulting row is nonzero, we normalize it to make the pivot equal to one and update the pivots array accordingly.
@@ 6.2 Multi-Modular Approach over the Rationals
In the commutative setting, multi-modular methods are typically used in an end-to-end fashion: a complete Gröbner basis is computed modulo different primes and only the final rational Gröbner basis is reconstructed. It has been shown recently [21] that this classical strategy fails in the noncommutative case because of the infinite nature of Gröbner bases in free algebras.
Nevertheless, while an end-to-end multi-modular computation is not viable in our setting, the linear algebra within the F4 algorithm can still be performed in a multi-modular fashion. Our implementation uses the multi-modular algorithm presented in [37, Alg. 7.6] for computing the reduced row echelon form of rational matrices efficiently. In this way, we can still circumvent expensive rational arithmetic and leverage fast modular computations. Moreover, we highlight that end-to-end methods lack a general and efficient test to verify the correctness of the reconstructed result and therefore yield correct results only with high probability [2, 6]. In contrast, our method provides a simple verification criterion (see line 10 of Alg. 1) and is thus guaranteed to return the correct result.
The multi-modular algorithm is summarized in Algorithm 1. In this description, we denote by $H(A)=\max_{i,j}\lvert A_{i,j}\rvert$ the height of an integer matrix $A$. For a correctness proof of Algorithm 1, we refer to [37, Alg. 7.6]. For further details on the reconstruction via Chinese remaindering and rational reconstruction, see, e.g., [15, Ch. 5].
@@
6:  discard any $R_{p}$ whose pivot column vector is not maximal  [^2] The pivot column vector of $R_{p}$ consists of the indices of all pivot columns of $R_{p}$ sorted in ascending order. The maximality of such a vector is determined w.r.t. the following ordering: first the dimensions are compared (shorter vectors are smaller) and vectors of the same dimension are ordered according to the reverse lexicographic order.among all $R_{p}$, $p\in P$
@@
To optimize Algorithm 1, we have implemented a technique described in [6]: During the very first modular computation in line 5, we construct a data structure called tracer, which records all rows of the matrix that reduce to zero. In all subsequent modular echelon form computations, we then apply this tracer to immediately zero out the corresponding rows, avoiding any additional computations. This optimization drastically speeds up the multi-modular algorithm, however the final output will only be correct with high probability. In particular, the result is guaranteed to be correct if the true reduced echelon form does not contain an entry whose denominator is divisible by the prime used to compute the tracer. This happens with high probability and the number of bad primes is finite. In our implementation, we use the Mersenne prime $p=2^{31}-1$ as the first prime for which the tracer is computed. This allows us to exploit the fast Mersenne modular operation discussed at the end of Section 6.1.
@@ 8 Experiments
We evaluate f4ncgb against the related tools from Section 2, focusing on GBNP, Letterplace, NCPoly, and operator_gb. We exclude bergman and Opal from the evaluation as they are no longer maintained and were already outperformed by earlier versions of Letterplace [29]. We also omit NCAlgebra as initial experiments indicated that it is several orders of magnitude slower than the other tools (NCAlgebra required almost 3 hours to solve braid3-11, while f4ncgb needed $0.1$ seconds). Due to licensing restrictions, we could not evaluate Magma. However, to offer a rough comparison, we include Magma timings from [32] on selected examples.
@@
Table 3 reports timings for the parallelized version of f4ncgb. For these tests, we use (larger) instances from the SymbolicData project and we limit our comparison to Letterplace, as all other tools performed significantly worse on the smaller examples. On these larger benchmarks, the single-threaded version of f4ncgb achieves speedups of up to $880\times$ over Letterplace and the parallelized version reaches up to $1500$\times$. All reported times are real (wall-clock) times.
@@
Example f4ncgb Letterplace Magma GBNP NCPoly operator_gb
lascala_neuh-10 $1.24$ $19.86$ $13.62$ $271.66$ $63.96$ $32.89$
serre-f4-15 $4.90$ $3.66$ $8.96$ $2262.04$ $4303.90$ $8404.25$
serre-ha11-15 $4.36$ $8.00$ $5.80$ $2348.60$ $6258.89$ $7793.97$
serre-eha112-13 $1.51$ $1.39$ $1.72$ $315.32$ $560.23$ $397.75$
4nilp5s-8 $3.24$ $24.48$ $9.54$ $1126.18$ $548.48$ $147.17$
braidXY-11 $15.86$ $87.97$ $4.20$ $120.55$ $7649.53$ $5196.03$
ug2-x1x2x3x4-5 $0.18$ $2.56$ $0.83$ $2.61$ $26.33$ $6.53$
serre-e6-15 $21.39$ $14.80$ $40.99$ $2968.67$ $8862.33$ TO
braid3-11 $0.10$ $1.36$ $0.64$ $120.55$ $107.48$ $55.04$
ufn3-10 $2.60$ $34.76$ $2.26$ $244.30$ $284.16$ $58.67$
ls3nilp-10 $0.21$ $0.53$ $1.98$ $21.76$ $10.27$ $13.32$
Table 2: Timings on selected benchmarks (in sec). For Magma, we report the timings from [32]. TO: $43\,200\text{\,}\mathrm{s}$\$
Example Letterplace f4ncgb 1 core 2 cores 4 cores 8 cores 16 cores 32 cores 4nilp5s-10 $1282.44$ $150.46$ $96.83$ $79.47$ $65.96$ $63.03$ $57.44$ braid3-16 18 953.1918\,953.19 $105.46$ $64.92$ $34.58$ $23.96$ $18.14$ $17.69$ braidX-18 TO $1977.27$ $1018.47$ $601.12$ $364.29$ $260.65$ $239.30$ braidXY-12 $1847.97$ $62.40$ $55.45$ $52.44$ $53.51$ $52.98$ $55.26$ holt_G3562h-17 TO 25 021.6225\,021.62 14 660.0414\,660.04 12 671.3812\,671.38 $8915.69$ $6824.36$ $5715.45$ lascala_neuh-13 $171.95$ $9.84$ $7.08$ $5.50$ $4.98$ $4.90$ $5.25$ lp1-15 24 166.8524\,166.85 $266.44$ $209.39$ $179.22$ $161.98$ $155.05$ $168.91$ lv2d10-100 TO $48.43$ $34.31$ $27.68$ $26.98$ $47.65$ $85.93$ malle_G12h-100 $4142.27$ $89.93$ $78.04$ $74.15$ $74.19$ $73.91$ $79.86$ serre-e6-17 $113.95$ $130.62$ $130.07$ $130.53$ $131.01$ $128.97$ $133.14$ serre-ha11-17 $99.09$ $38.18$ $32.79$ $31.66$ $30.77$ $30.84$ $31.87$
Table 3: Timings on selected benchmarks (in sec). TO: $43\,200\text{\,}\mathrm{s}$\$

### v2

@@ Abstract
We present f4ncgb, a new open-source C++ library for Gröbner basis computations in free algebras, which transfers recent advancements in commutative Gröbner basis software to the noncommutative setting. As our experiments show, f4ncgb establishes a new state of the art for noncommutative Gröbner basis computations. We also discuss implementation details and design choices.
@@ 1 Introduction
In this work, we present f4ncgb, a new open-source C++ library for Gröbner basis computations in free algebras, distributed under the license MIT and available at https://gitlab.sai.jku.at/f4ncgb/f4ncgb. The library supports noncommutative polynomial systems with coefficients in prime fields of characteristic $2^{31}$ and in the field of rational numbers. To the best of our knowledge, f4ncgb is the first open-source tool to port the newest advancements in Gröbner basis computations from the commutative setting to the noncommutative world. Our experiments in Section 8 show that f4ncgb establishes a new state of the art for noncommutative Gröbner basis computations among freely available tools.
The foundation of our implementation are specialized data structures for representing and storing noncommutative monomials and polynomials. We abstract these data types into a store structure, using a contiguous memory area for storing elements and a hash map for efficient lookup of known entries. We describe this structure in more detail in Section 5.
@@
To efficiently support this, we have implemented a specialized Gaussian elimination algorithm tailored for sparse matrices, based on ideas from the commutative case [34]. Our implementation makes use of modern multi-core hardware by parallelizing matrix reductions, resulting in significant overall speedups. For inputs with rational coefficients, we employ a multi-modular strategy: computations are carried out modulo different primes, followed by a reconstruction step to recover the rational solution. We provide an overview of the noncommutative F4 algorithm in Section 4 and discuss our linear algebra implementation in Section 6.
In addition to computing Gröbner bases, f4ncgb offers proof logging capabilities in the form of cofactor representations: for each polynomial in the output basis, a representation in terms of the input can be computed, see Section 7. These representations are not only of interest for verifying the correctness of the output, but are also useful in applications [23].
@@ 2 Related Software
Notable examples of tools relying on Buchberger’s algorithm include the Mathematica library NCAlgebra [18], the ApCoCoA package NCPoly [29], and the GAP package GBNP [11]. Other tools that also fall into this category are the Standard Lisp implementation bergman [3] and the C++ tool Opal [17], which both seem no longer actively maintained.
An important exception is the subscription-based computer algebra system Magma [8], which is one of the few tools to implement a noncommutative version of the F4 algorithm. In addition, Magma also features dedicated data structures and highly optimized subroutines (e.g., for the linear algebra), making it arguably the state of the art for computing noncommutative Gröbner bases. However, a significant drawback is its proprietary nature, which limits its availability to the community.
For homogeneous ideals over prime fields, the open-source system Macaulay2 [16] offers a decent implementation of the F4 algorithm. For all other ideals, however, it falls back to Buchberger’s algorithm.
To the best of our knowledge, the only open-source implementation of the noncommutative F4 algorithm over the field of rational numbers is the SageMath package operator_gb [5], developed by one of the authors. While operator_gb already implements the noncommutative F4 algorithm, it still relies on relatively simple data structures for representing and manipulating noncommutative objects. Additionally, the package is primarily written in interpreted Python, with selected components (such as the linear algebra) in compiled Cython, which limits performance. As a result, it cannot match the efficiency of the highly optimized and fully compiled C code in Magma. The new package f4ncgb can be viewed as a significant advancement of operator_gb, offering both improved algorithmic design and a more efficient implementation.
@@ 3 Preliminaries – Noncommutative Gröbner Bases
For a set of polynomials $F\subseteq K\langle X\rangle$, we let $(F)$ be the (two-sided) ideal generated by $F$, which consists of all two-sided linear combinations of the elements in $F$ with polynomials as coefficients, i.e., $(F)=\left\{\sum_{i=1}^{k}p_{i}f_{i}q_{i}\mid f_{i}\in F,\;p_{i},q_{i}\in K\langle X\rangle,\;k\in\mathbb{N}\right\}$. A set $F$ is a generating set of an ideal $I\subseteq K\langle X\rangle$ if $I=(F)$. Recall that if $|X|>1$, then there exist ideals in $K\langle X\rangle$ that do not possess a finite generating set. A classical example is the ideal $(\{xy^{n}x\mid n\in\mathbb{N}\})\subseteq K\langle x,y\rangle$.
@@
A very general class of monomial orderings are weighted block orderings of the following form: Fix $d\in\mathbb{N}_{>0}$ and assign to each variable $x_{i}\in X$ a nonzero weight vector $w_{i}\in\mathbb{R}_{\geq 0}^{d}$. Define the weight $w(m)$ of a monomial $m=x_{i_{1}}\dots x_{i_{k}}\in\langle X\rangle$ as $w(m)=\sum_{j=1}^{k}w_{i_{j}}\in\mathbb{R}_{\geq 0}^{d}$. The weighted block ordering $\prec_{\omega}$ w.r.t. the weight matrix $\omega=(w_{1},\dots,w_{n})$ is given by:
@@
At the moment, our implementation supports weighted block orderings where each weight vector is a unit vector $(0,\dots,0,1,0,\dots,0)$. This is sufficient to represent classical orderings such as degree-lexicographic orderings. It also enables the use of elimination orderings [21, Def. 2.4.15], which are crucial, e.g., for computing elimination ideals [21, Thm. 2.4.43].
@@
We note that all standard characterizations of Gröbner bases from the commutative setting carry over naturally to the noncommutative case; see, e.g., [21, Thm. 2.4.37]. In particular, Buchberger’s criterion [13, Thm. 2.6.6], which provides an effective test for whether a set of commutative polynomials forms a Gröbner basis, has a noncommutative analogue in Bergman’s diamond lemma [4, Thm. 1.2].
@@
Analogous to Buchberger’s criterion, Bergman’s diamond lemma is based on a notion of (noncommutative) polynomial reduction, which is (basically) identical to the reduction used in the commutative case: $f$ can be reduced by $g$ if $\operatorname{lm}(g)$ divides a monomial in $f$ and the result of the reduction is $f-\lambda vgw$ with $\lambda\in K$, $v,w\in\langle X\rangle$ chosen so that the targeted monomial cancels. For a formal definition, see, e.g., [21, Sec. 2.4.2].
@@
Theorem 3.1 provides an effective criterion to test whether a given set forms a Gröbner basis. It also directly gives rise to a completion procedure for constructing a Gröbner basis from an arbitrary generating set – a noncommutative analogue of Buchberger’s algorithm; see [21, Alg. 1] for a detailed description.
@@ 4 Noncommutative F4 Algorithm
Our library f4ncgb implements the noncommutative F4 algorithm for computing Gröbner bases in free algebras. Below, we provide a brief overview of the algorithm; for a detailed exposition, see [40, Sec. 5.4]. For those familiar with the commutative algorithm, we note that the noncommutative version follows the same high-level structure.
After initializing the prospective Gröbner basis with the input polynomials and computing all ambiguities between them, the F4 algorithm enters a loop consisting of the following main steps:
@@
    Basis update: Translate the rows of the reduced matrix back into polynomials and add all those polynomials to the basis whose leading monomial is not divisible by an element of the current basis. Then, form all new ambiguities. Optionally, one can apply the noncommutative variant of the Gebauer-Möller criteria [40, Thm. 4.2.22] to discard redundant ambiguities. If ambiguities remain, continue with the next iteration.
@@ 5 Basic Data Structures – Monomials and Polynomials
To store monomials (and also polynomials), we have designed a common base store structure that is extended into monomial store and polynomial store structures using the Curiously Recurring Template Pattern [12]. The store implements a bump-pointer allocator combined with an open-addressing hash map (using Boost’s unordered_flat_map) for efficient lookups of known entries.
@@
[…]
@@ 6.1 Gaussian Elimination in Prime Fields
The forward elimination phase begins with a preprocessing step in which we permute the rows of $A$ such that, in each row, the first nonzero entry appears directly below or to the left of the first nonzero entry in the row above. This reordering yields a matrix with an approximately triangular structure, as illustrated in Figure 2. We note that this preprocessing step is crucial for the efficiency of the algorithm.
currently processed row(dense buffer)already knownpivot rows (sparse)
@@
Once the buffer is fully reduced, we copy its content back into the sparse row $r$. If the resulting row is nonzero, we normalize it to make the pivot equal to $1$ and update the pivots array accordingly.
@@ 6.2 Multi-Modular Approach over the Rationals
In the commutative setting, multi-modular methods are typically used in an end-to-end fashion: a complete Gröbner basis is computed modulo different primes and only the final rational Gröbner basis is reconstructed. It has been shown [22] that this classical strategy fails in the noncommutative case because of the infinite nature of Gröbner bases in free algebras.
Nevertheless, while an end-to-end multi-modular computation is not viable in our setting, the linear algebra within the F4 algorithm can still be performed in a multi-modular fashion. Our implementation uses the multi-modular algorithm presented in [38, Alg. 7.6] for computing the reduced row echelon form of rational matrices efficiently. In this way, we can still circumvent expensive rational arithmetic and leverage fast modular computations. Moreover, we highlight that end-to-end methods lack a general and efficient test to verify the correctness of the reconstructed result and therefore yield correct results only with high probability [2, 6]. In contrast, our method provides a simple verification criterion (see line 10 of Alg. 1) and is thus guaranteed to return the correct result.
The multi-modular algorithm is summarized in Algorithm 1. In this description, we denote by $H(A)=\max_{i,j}\lvert A_{i,j}\rvert$ the height of an integer matrix $A$. For a correctness proof of Algorithm 1, we refer to [38, Alg. 7.6]. For further details on the reconstruction via Chinese remaindering and rational reconstruction, see, e.g., [15, Ch. 5].
@@
6:  discard any $R_{p}$ whose pivot column vector is not maximal  [^2] The pivot column vector of $R_{p}$ consists of the indices of all pivot columns of $R_{p}$ sorted in ascending order. The maximality of such a vector is determined as follows: first the dimensions are compared (shorter vectors are smaller) and vectors of the same dimension are ordered according to the reverse lexicographic order.among all $R_{p}$, $p\in P$
@@
To optimize Algorithm 1, we have implemented a technique described in [6]: During the very first modular computation in line 5, we construct a data structure called tracer, which records all rows of the matrix that reduce to zero. In all subsequent modular echelon form computations, we then apply this tracer to immediately zero out the corresponding rows, avoiding any additional computations. This optimization drastically speeds up the multi-modular algorithm, however the final output will only be correct with high probability. In particular, the result is guaranteed to be correct if the true reduced echelon form does not contain an entry whose denominator is divisible by the prime used to compute the tracer. This happens with high probability and the number of bad primes is finite. In our implementation, we use the Mersenne prime $p=2^{31}-1$ as the first prime for which the tracer is computed. This allows us to exploit the fast Mersenne modular operation discussed at the end of Section 6.1. The tracer can be disabled to ensure correctness of the output.
@@ 8 Experiments
We evaluate f4ncgb against the related tools from Section 2, focusing on GBNP, Letterplace, Macaulay2, NCPoly, and operator_gb. We exclude bergman and Opal from the evaluation as they are no longer maintained and were already outperformed by earlier versions of Letterplace [30]. We also omit NCAlgebra as initial experiments indicated that it is several orders of magnitude slower than the other tools (NCAlgebra required almost 3 hours to solve braid3-11, while f4ncgb needed $0.1$ seconds). Due to licensing restrictions, we could not evaluate Magma. However, to offer a rough comparison, we include Magma timings from [33].
@@
Table 3 reports timings for the parallelized version of f4ncgb. For these tests, we use (larger) instances from the SymbolicData project and we limit our comparison to Letterplace and Macaulay2, as all other tools performed significantly worse on the smaller examples. On these larger benchmarks, the single-threaded version of f4ncgb achieves speedups of up to $880\times$ over Letterplace and up to $510\times$ over Macaulay2. The parallelized version reaches speedups of up to $1500$\times$ and $920$\times$, respectively. All reported times are real (wall-clock) times.
@@
Example f4ncgb Letterplace Macaulay2 Magma GBNP NCPoly operator_gb
lascala_neuh-10 $1.24$ $19.86$ $7.57$ $13.62$ $271.66$ $63.96$ $32.89$
serre-f4-15 $4.90$ $3.66$ $5.63$ $8.96$ $2262.04$ $4303.90$ $8404.25$
serre-ha11-15 $4.36$ $8.00$ $12.05$ $5.80$ $2348.60$ $6258.89$ $7793.97$
serre-eha112-13 $1.51$ $1.39$ $3.91$ $1.72$ $315.32$ $560.23$ $397.75$
4nilp5s-8 $3.24$ $24.48$ $18.91$ $9.54$ $1126.18$ $548.48$ $147.17$
braidXY-11 $15.86$ $87.97$ $1305.72$ $4.20$ $120.55$ $7649.53$ $5196.03$
ug2-x1x2x3x4-5 $0.18$ $2.56$ TO $0.83$ $2.61$ $26.33$ $6.53$
serre-e6-15 $21.39$ $14.80$ $14.51$ $40.99$ $2968.67$ $8862.33$ TO
braid3-11 $0.10$ $1.36$ $5.20$ $0.64$ $120.55$ $107.48$ $55.04$
ufn3-10 $2.60$ $34.76$ $6.05$ $2.26$ $244.30$ $284.16$ $58.67$
ls3nilp-10 $0.21$ $0.53$ $2.17$ $1.98$ $21.76$ $10.27$ $13.32$
Table 2: Timings (in sec) on benchmarks from [33]. For Magma, we report the timings from [33]. TO: $43\,200\text{\,}\mathrm{s}$\$
Example Letterplace Macaulay2 f4ncgb 1 core 4 cores 8 cores 16 cores 32 cores 4nilp5s-10 $1282.44$ $875.09$ $150.46$ $79.47$ $65.96$ $63.03$ $57.44$ braid3-16 18 953.1918\,953.19 14 291.3714\,291.37 $105.46$ $34.58$ $23.96$ $18.14$ $17.69$ braidX-18 TO TO $1977.27$ $601.12$ $364.29$ $260.65$ $239.30$ braidXY-12 $1847.97$ 18 887.6118\,887.61 $62.40$ $52.44$ $53.51$ $52.98$ $55.26$ holt_G3562h-17 TO TO 25 021.6225\,021.62 12 671.3812\,671.38 $8915.69$ $6824.36$ $5715.45$ lascala_neuh-13 $171.95$ $37.94$ $9.84$ $5.50$ $4.98$ $4.90$ $5.25$ lp1-15 24 166.8524\,166.85 33 923.1733\,923.17 $266.44$ $179.22$ $161.98$ $155.05$ $168.91$ lv2d10-100 TO 24 930.2524\,930.25 $48.43$ $27.68$ $26.98$ $47.65$ $85.93$ malle_G12h-100 $4142.27$ $163.38$ $89.93$ $74.15$ $74.19$ $73.91$ $79.86$ serre-e6-17 $113.95$ $84.99$ $130.62$ $130.53$ $131.01$ $128.97$ $133.14$ serre-ha11-17 $99.09$ $108.41$ $38.18$ $31.66$ $30.77$ $30.84$ $31.87$
Table 3: Timings on larger benchmarks (in sec). TO: $43\,200\text{\,}\mathrm{s}$\$